OpenStudy (anonymous):

A student increases the temperature of a 300 cm3 balloon from 30°C to 60°C. What will the new volume of the balloon be? (Remember to be careful with the units.) 273 cm3 300 cm3 600 cm3 329 cm3

OpenStudy (anonymous):

is it 273?

OpenStudy (anonymous):

@Riddellikins

OpenStudy (anonymous):

PV = NrT We can assume pressure stays the same (it equalises as the balloon gets bigger) and the number of molecules along with the universal gas constant are the same to bring us to the same equation as before: V = kT Use the beginning values to find k 300 = k * 30 k = 10 Use the original equation and k to find the missing value V = kT V = 10 * 60 V = 600 Therefore the volume is 600cm3

OpenStudy (anonymous):

thank you so much, can you check this one too? You place air in a sealed can at standard temperature and pressure (STP). You double its absolute temperature (K) while leaving the volume constant. What happens to the pressure? It decreases by one-half.***** It stays the same. It doubles. It increases to four times the original pressure.

OpenStudy (anonymous):

PV = NrT Amount of molecules, Ugas constant and volume are constant: P = kT Therefore if the temperature doubles, the pressure will too. Think about it logically though. If you increase the temperature of a sealed can then the pressure will rise, because eventually it will explode.